Two small conducting and identical spheres A and B have charges -25 nC and +15 nC, respectively. They are separated by a distance of 0.02 m.


1
Expert's answer
2022-02-15T12:53:02-0500

Two small conducting and identical spheres A and B have charges -25 nC and +15nC, respectively. They are separated by a distance of 0.02m. What is the magnitude of the electric force between the two spheres?

Solution:

Given:

"q_{A}=-25\ nC=-25\times10^{-9}\ nC, q_B=15\ nC=15\times10^{-9}\ nC\\
d=0.02 \ m"

Electric force

"F=\frac{1}{4\pi\epsilon} \frac{q_A q_B}{d^2}=9\times10^9\times \frac{-25\times10^{-9}\times15\times10^{-9}}{0.02^2}=8.4375\times10^{-3} \ N"